OpenStudy (mockingbird):

calculus---- plane areas in polar coordinates:

OpenStudy (mockingbird):

Find the area in the first quadrant under the curve y^2 = x^2/ x-1 between the minimum ordinate and the line x=3

OpenStudy (mockingbird):

y= 3/2

OpenStudy (mockingbird):

intersection

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\int\limits_{\theta1}^{\theta2}f(\theta)^2 d \theta\] This will help find the area once you find the endpoints and the function in polar.

OpenStudy (anonymous):

Some other helpful equations: \[x=rcos (\theta), y=rsin(\theta)\] The function f(x) can be converted to\[f(\theta)\] from this.